# Eco-evolutionary dynamics of cooperative antimicrobial resistance in a population of fluctuating volume and size

### v1

@@ 4 Dynamic environment
Environments are often dynamic and endlessly change through time. Having characterised how the system evolves when the environment is static, we now look to include the biologically-motivated environmental changes driven by the binary switching of the carrying capacity (5). These environmental switches can produce population bottlenecks, where the community size shrinks significantly and is subject to strong demographic fluctuations, hence moulding the fixation/coexistence properties of the population.
@@
When the population is large enough for demographic fluctuations to be negligible, and randomness only arises from the time-fluctuating environment, via the random switches of the carrying capacity (5), the dynamics is well described by a so-called piecewise deterministic Markov process (PDMP) [Dav84]. In the PDMP, the total population dynamics between each environmental switch is deterministic and given by equation (7), with $K_{0}$ now replaced by $K_{\pm}$ when $\xi=\pm 1$. Since equations (7) and (8) are decoupled, the PDMP does not affect the mean-field dynamics of the population composition $x$, that is still given by (8). Here, we have PDMP for the population size $N$ ($N$-PDMP) that is therefore defined by
@@
where the time-varying carrying capacity $K(t)\in\{K_{-},K_{+}\}$ is given by equation (5). Sample paths of this $N$-PDMP are shown as solid black lines in figures 1(c) and 3(e)-(h). These realisations illustrate how $N(t)$ tracks the switching carrying capacity $K(t)$ for low and intermediate/low switching rates $\nu$, while it fluctuates about an effective value ${\cal K}$ for very high switching rates, see figure 3(h). The population composition $x(t)$ always tends towards the coexistence equilibrium $x\to x_{\text{th}}$; see figure 3(a)-(d). The dynamics of $K(t)$ leads us to identify two limiting regimes. When environmental variability is slow, $\nu\rightarrow 0$, the environmental state switches very rarely, and $N$ thus essentially coincides with $K(0)$ giving $N=K_{\pm}$ with probability $(1\pm\delta)/2$. In this slow EV regime, the population dynamics is the same as in static environments with a constant carrying capacity set by the initial condition. When the environment switches very frequently, $\nu\rightarrow\infty$, $N$ cannot keep track of the changes of $K(t)$, which approaches an effective value $K(t)\to{\cal K}$ obtained by self-averaging the environmental noise, yielding $\mathcal{K}=1/\langle 1/K(t)\rangle=2K_{+}K_{-}/\left[(1-\delta)K_{+}+(1+\delta)K_{-}\right]$ [WFM17, WFM18, WM20, SMM21, TWAM20, TWMA23, HNARM23a, AHNRM23]. Under very high switching rate, the population size thus approaches this effective constant carrying capacity $N\approx{\cal K}$, see figure 3(h).
The (marginal) stationary probability density function (pdf) of the $N$-PDMP (ignoring demographic fluctuations) for fixed $\{\nu,\delta\}$ is [Hor84, HL06, Ben06, RDL11, WFM17, WFM18, TWAM20, AHNRM23]
@@
where $\mathcal{Z}$ is a normalisation constant and $N\in\left[K_{-},K_{+}\right]$ is a continuous variable. Despite ignoring demographic fluctuations, the above equation provides a useful approximation of the actual quasi-stationary population size distribution, and captures well the underlying total population dynamics [WFM17, WFM18, WM20, TWAM20, AHNRM23]. When $\nu\ll 1$ the pdf is bimodal and sharply peaked at $N=K_{-}$ and $N=K_{+}$. The probability density $p_{\nu,\delta}(N)$ begins to flatten out as $\nu\sim 1$. When the switching rate $\nu$ is increased further, the pdf becomes unimodal and sharply peaked about a value approaching $\mathcal{K}$, and eventually $N\approx\mathcal{K}$, when $\nu\gg 1$. The actual $N$ quasi-stationary distribution shows slightly broader peaks, due to demographic noise, but is otherwise well described by the PDMP approximation [WFM17, WFM18, TWAM20, TWMA23]. In particular, the average of $N$ over the $N$-PDMP pdf given by
@@
is an accurate approximation of $\langle N\rangle$, the actual stationary mean population size [WFM17, WFM18, TWAM20, AHNRM23], with $\langle N\rangle\approx\langle N\rangle_{\nu,\delta}\approx[(1+\delta)K_{+}+(1-\delta)K_{-}]/2$ when $\nu\ll 1$ and $\langle N\rangle\approx\langle N\rangle_{\nu,\delta}\approx\mathcal{K}$ when $\nu\gg 1$. As shown in references [WFM17, WFM18, TWAM20, AHNRM23], equation (18) correctly captures that, at fixed $\delta$, the average population size $\langle N\rangle$ is a decreasing function of $\nu$.
Figure 3: Eco-evolutionary dynamics sample paths. (a)-(d) Examples of $R$ fraction sample paths in silico (blue lines) for four concentration cooperation thresholds $x_{\text{th}}$ (dashed black lines), three average switching rates $\nu$, and two environmental biases $\delta$, for $s=0.1$, $a=0.25$, $K_{-}=100$, and $K_{+}=1000$, with initial conditions $N_{R}(t=0)=x_{\text{th}}K(t=0)$ and $N_{S}(t=0)=\left(1-x_{\text{th}}\right)K(t=0)$. Shaded and white areas in panels (b)-(c) encode periods of abundance ($K(t)=K_{+}$) and scarcity ($K_{-}$), respectively. Note the larger amplitude of demographic fluctuations for the latter since $N\rightarrow K_{-}\ll K_{+}$. (e)-(h) Same example paths as in the corresponding panels (a)-(d) for the population size $N$ (green lines); black lines show example path for the carrying capacity $K$. The very high environmental switching rate $\nu$ in panel (h) provides the effectively constant carrying capacity $\mathcal{K}$ (black dashed line); see section 4.1. (i)-(l) Same example paths as in the previous panels for the number of $R$ (blue lines) and $S$ (red lines). Dashed lines show the corresponding (meta-)stable coexistence equilibrium $N_{R}=x_{\text{th}}K$ (blue) and $N_{S}=(1-x_{\text{th}})K$ (red); see section 3. The demographic noise (DN) and the low value of $x_{\text{th}}=0.1$ in panels (a),(e),(i) lead to fast extinction of $R$, whereas the high threshold $x_{\text{th}}=0.9$ and DN in panels (d),(h),(l) lead to an early fixation of $R$ and extinction of $S$.
@@ 4.3 Theory for the fixation-coexistence diagrams in fluctuating environments
In this section we develop the analytical methods for fluctuating environments allowing us to predict when, in our simulations, one of the strains will fixate or go extinct, and when both strains will coexist for long-periods. The theory that we have devised and discuss here is used to analytically reproduce, see figure 4(e)-(h), the fixation-coexistence diagrams of figure 4(a)-(d) obtained from extensive computer simulations.
@@
We first define long-lived coexistence in fluctuating environments by direct extension of what has been discussed in section 3.2. For given EV defined by the environmental parameters $\{\nu,\delta\}$, there is long-lived coexistence of the strain $R$ and $S$ if the fixation time, or equivalently, the extinction time, here denoted by $\tau$, exceeds $2\langle N\rangle$ [HNARM23a, AHNRM23]. Note that the mean fixation/extinction time, denoted by $\langle\tau\rangle$, coincides with the mean coexistence time, as in the environment-static case of section 3.2. We will thus say that, for given $\{\nu,\delta\}$, there is long-lived coexistence in the population of fluctuating size and volume, if $\tau>2\langle N\rangle\approx 2\langle N\rangle_{\nu,\delta}$. The rationale for this definition is the same as in the case of static environment, see section 3.2 and [CRF09, HMT11, RMF07, HNARM23a, AHNRM23], with the difference that the constant population is now replaced by the average stationary population size $\langle N\rangle$ that we analytically approximate by $\langle N\rangle_{\nu,\delta}$ obtained from the $N$-PDMP according to equation (18). Therefore, we will say that, for given environmental parameters $\{\nu,\delta\}$, there is fast fixation or, equivalently, early extinction of a strain, if $R$ or $S$ fixates in a time $\tau$ that is less than $2\langle N\rangle\approx 2\langle N\rangle_{\nu,\delta}$.
@@
Secondly, we distinguish two fluctuating environment regimes in which the fixation and long-lived coexistence probability are evaluated by distinct suitable approaches. We refer to these different behaviours as the quenched (Q) and annealed (A) regimes. Dynamic environments in the regime Q are so slow, $\nu\ll 1$, and/or so biased, $\delta\approx\pm 1$, that the environment is unlikely to experience any switch by time $2\langle N\rangle_{\nu,\delta}$ regardless of the initial condition; whereas, in the same time, environments in regime A experience at least one environmental switch. Similar Q and A environmental regimes have been identified elsewhere, see e.g. [MTAS23, Mob23] [^3] It is worth noting that the annealed regime often corresponds to the limit $\nu\to\infty$ where the environmental noise $\xi$ self averages [MTAS23, Mob23, WFM17, WFM18, TWAM20, TWMA23]. Here, the case of intermediate switching with a finite number of switches prior to fixation/extinction is qualitatively similar to the case $\nu\to\infty$, and for simplicity we here refer to it as the annealed regime A.. The border between Q and A regimes is determined by
@@
which splits the fixation-coexistence $\nu-\delta$ diagrams of figure 4 in two regions separated by green/yellow lines: where $1/\left(\nu(1-|\delta|)\right)>2\langle N\rangle_{\nu,\delta}$, the average time for an environmental switch to occur exceeds the reference time for long-lived coexistence. This means that in this region $K$ is unlikely to experience any switch by time $2\langle N\rangle_{\nu,\delta}$, and $2\left(\nu(1-|\delta|)\right)<1/\langle N\rangle_{\nu,\delta}$ hence delineates the Q regime. Here, the A regime is thus defined by $1/\left(\nu(1-|\delta|)\right)<2\langle N\rangle_{\nu,\delta}$, where, on average, $K$ switches at least once by time $2\langle N\rangle_{\nu,\delta}$.
Quenched environmental regime. In regime Q (on the left of green/yellow lines of figure 4), $K(0)=K_{+}$ with probability $(1+\delta)/2$ or $K(0)=K_{-}$ with probability $(1-\delta)/2$, and the carrying capacity remains at initial value, $K(t)=K(0)$, at least until time $2\langle N\rangle_{\nu,\delta}$. Hence, the joint probability of $R$ fixating and that fixation/extinction occurs in a time $\tau<2\langle N\rangle_{\nu,\delta}$ (fast fixation) in the regime Q, denoted by $\Phi^{\text{Q}}$, is
@@
is the probability of long-lived coexistence in the Q regime, starting (and staying) at a total population $N$, with MCT $\langle\tau_{N}\rangle$, where $\eta_{N}^{Q}$ is defined as the environmental-dynamic counterpart of equation (14).
Annealed environmental regime. In regime A (on the right of green/yellow lines of figure 4), the environment switches at least once prior to time $2\langle N\rangle_{\nu,\delta}$ regardless of the initial condition. In this regime, $N$ therefore experiences a broad range of values in sampling its quasi-stationary distribution, here approximated by $p_{\nu,\delta}(N)$ of equation (17), before a fixation/extinction event or long-lived coexistence occurs. In the vein of [WFM17, WFM18], we thus assume that when a fixation event occurs, it arises in a population of size $N$ ponderated by the probabilistic weight $p_{\nu,\delta}(N)$. Following [MA10, AM10], we know that the MCT $\langle\tau_{N}\rangle$ in a population of constant size $N$ is the inverse of the flux towards the absorbing boundaries, and the fixation probability $\phi_{N}$ is given by the relative flux into the absorbing state $N_{R}=N$ [AM10]. Hence, $\phi_{N}/\langle\tau_{N}\rangle$ is defined as the “rate to $R$ fixation” in a population of size $N$, and similarly we define $(1-\phi_{N})/\langle\tau_{N}\rangle$ as the “rate to $R$ extinction” or, equivalently, as the “rate to $S$ fixation”. In regime A, we assume that any fixation/extinction events occur while $N$ is at quasi-stationarity and, by analogy with the $N$-constant case, experiences a rate to $R$ fixation/extinction $\phi_{N}/\langle\tau_{N}\rangle$ and $(1-\phi_{N})/\langle\tau_{N}\rangle$, respectively; but we now average these over $p_{\nu,\delta}(N)$ to account for the varying population size in fluctuating environments. We thus obtain the effective rates of $R$ fixation/extinction in regime A, respectively given by
@@
Since $\langle\tau_{K_{+}}\rangle\gg\langle\tau_{K_{-}}\rangle$, the main contributions to $\mathcal{T}_{F/E}$, and hence to fixation/extinction, arise from $N\approx K_{-}$. With $\mathcal{T}_{F/E}$, we obtain the $R$ fixation probability and MCT, $\phi^{\text{A}}$ and $\langle\tau^{\text{A}}\rangle$, respectively, in formal analogy with the environmental-static case [MA10, AM10]:
@@
Moreover, for a dynamic environment of parameters $\{\nu,\delta\}$, following the reasoning leading to equation (14), the long-lived coexistence probability in regime A is
@@
Crossover regime and general results. Having obtained the fixation and coexistence probabilities in regimes Q and A, we can suitably superpose their expressions to obtain predictions applicable in the crossover regime (about the green/yellow line in figure (4)), as well as in regimes Q and A. This provides us with general results, that are valid for then entire range of environmental parameters $\{\nu,\delta\}$. Since the probability that no switches occur by time $2\langle N\rangle_{\nu,\delta}$ is
@@
and the probability that at least one switch has occurred by $2\langle N\rangle_{\nu,\delta}$ is $1-\Pi$, the overall joint probability of $R$ fixation and fast fixation is
@@
The fixation-coexistence diagrams in figure 4(a)-(d) show simulation results for the probability of fast $R$ fixation (blue), fast $S$ fixation (red), or long-lived coexistence (black), see appendix section A, for different values of the concentration cooperation threshold $x_{\text{th}}$ when $N$ is at quasi-stationarity. The main features of these diagrams can be understood in terms of the analysis carried in section 3.2 and by referring to the critical cooperation threshold value $x_{\text{th}}^{*}\approx[\ln\left(1-s\right)]/[\ln\left(1-a\right)]$  [^4] In principle, a suitable environment-dynamic counterpart of $x_{\text{th}}^{*}$, given by equation (11), in a population of average size $\langle N\rangle_{\nu,\delta}$ is $x_{\text{th},\nu,\delta}^{*}\equiv\frac{\ln\left(1-s\right)}{\ln\left(1-a\right)}+\frac{1}{\langle N\rangle_{\nu,\delta}}\frac{\ln\left(\frac{s(1-a)}{a-s}\right)}{\ln\left(1-a\right)}$. However, when $\langle N\rangle\gg 1$ and $a$ is not too close to 1, as in our examples, $x_{\text{th},\nu,\delta}^{*}$ is well approximated by the leading contribution to $x_{\text{th}}^{*}$, and hence we can consider $x_{\text{th},\nu,\delta}^{*}\approx x_{\text{th}}^{*}\approx\frac{\ln\left(1-s\right)}{\ln\left(1-a\right)}$.: when $x_{\text{th}}<x_{\text{th}}^{*}$ (and $x_{\text{th}}\lesssim 1/\sqrt{\langle N\rangle_{\nu,\delta}}$), the most likely outcome is the fast fixation of $S$ along with early extinction of $R$, while fast fixation of $R$ (early extinction of $S$) is the most probable outcome when $x_{\text{th}}>x_{\text{th}}^{*}$ (and $1-x_{\text{th}}\lesssim 1/\sqrt{\langle N\rangle_{\nu,\delta}}$), and long-lived coexistence of $R$ and $S$ is expected when $x_{\text{th}}$ is close to $x_{\text{th}}^{*}$ (and/or when ${\rm min}(x_{\text{th}},1-x_{\text{th}})\gg 1/\sqrt{\langle N\rangle_{\nu,\delta}}$). For the parameters of figure 4, $x_{\text{th}}^{*}\approx 0.366$, and we indeed notice that most of figure 4(a) appears in red, panels (c) and (d) of figure 4 are mostly coded in blue, and most of figure 4(b) appears in black.

### v2

@@ 4 Dynamic environment
Environments are often dynamic and endlessly change through time. Having characterised how the system evolves when the environment is static, we now look to include the biologically-motivated environmental changes driven by the binary switching of the carrying capacity (5). These environmental switches can produce population bottlenecks, where the community size shrinks significantly and is subject to strong demographic fluctuations, hence moulding the fixation/coexistence properties of the population. In this section, we derive general analytical and numerical methods to assess the impact of environmental variability (EV) on population dynamics. In particular, we show that, in a dynamic environment, fast fixation of either strain is also possible when $x_{\text{th}}\approx x_{\text{th}}^{*}$, and long-lived coexistence can also occur when $x_{\text{th}}\approx 0$ or $1$, see section 4.2.
@@
Here we show how the EV shapes the total microbial population size irrespective of composition dynamics. On the one hand we show that slow-switching environments (with small mean switching frequency $\nu$) yield a bimodal distribution of microbial populations, narrowly peaked about either the low ($K_{-}$) or high ($K_{+}$) carrying capacity. On the other hand, fast-switching EV (high $\nu$) leads to a unimodal distribution of the total population size peaked at an effective, intermediate carrying capacity $\mathcal{K}$. The intermediate-switching regime is characterised by a mixed behaviour and a wider variability of the population distribution.
When the population is large enough for demographic fluctuations to be negligible, and randomness only arises from the time-fluctuating environment, via the random switches of the carrying capacity (5), the dynamics is well described by a so-called piecewise deterministic Markov process (PDMP) [94]. In the PDMP, the total population dynamics between each environmental switch is deterministic and given by equation (7), with $K_{0}$ now replaced by $K_{\pm}$ when $\xi=\pm 1$. Since equations (7) and (8) are decoupled, the PDMP does not affect the mean-field dynamics of the population composition $x$ (and vice versa), which is still given by (8). Here, the PDMP for the population size $N$ ($N$-PDMP) is therefore defined by
@@
where the time-varying carrying capacity $K(t)\in\{K_{-},K_{+}\}$ is given by equation (5). Sample paths of this $N$-PDMP are shown as solid black lines in figures 1(c) and 3(e)-(h). These realisations illustrate how $N(t)$ tracks the switching carrying capacity $K(t)$ for low and intermediate/low switching rates $\nu$, while it fluctuates about an effective value ${\cal K}$ for very high switching rates, see figure 3(h). The population composition $x(t)$ always tends towards the coexistence equilibrium $x\to x_{\text{th}}$; see figure 3(a)-(d). The dynamics of $K(t)$ leads us to identify two limiting regimes. When environmental variability is slow, $\nu\rightarrow 0$, the environmental state switches very rarely, and $N$ thus essentially coincides with $K(0)$ giving $N=K_{\pm}$ with probability $(1\pm\delta)/2$. In this slow EV regime, the population dynamics is the same as in static environments with a constant carrying capacity set by the initial condition. When the environment switches very frequently, $\nu\rightarrow\infty$, $N$ cannot keep track of the changes of $K(t)$, which approaches an effective value $K(t)\to{\cal K}$ obtained by self-averaging the environmental noise, yielding $\mathcal{K}=1/\langle 1/K(t)\rangle=2K_{+}K_{-}/\left[(1-\delta)K_{+}+(1+\delta)K_{-}\right]$ [11, 12, 82, 46, 40, 41, 32, 13]. Under very high switching rate, the population size thus approaches this effective constant carrying capacity $N\approx{\cal K}$, see figure 3(h).
The (marginal) stationary probability density function (pdf) of the $N$-PDMP (ignoring demographic fluctuations) for fixed $\{\nu,\delta\}$ is [95, 80, 79, 81, 11, 12, 40, 13]
@@
where $\mathcal{Z}$ is a normalisation constant and $N\in\left[K_{-},K_{+}\right]$ is a continuous variable. Despite ignoring demographic fluctuations, the above equation provides a useful approximation of the actual quasi-stationary population size distribution, and captures well the underlying total population dynamics [11, 12, 82, 40, 13]. When $\nu\ll 1$ the pdf is bimodal and sharply peaked at $N=K_{-}$ and $N=K_{+}$. The probability density $p_{\nu,\delta}(N)$ begins to flatten out as $\nu\sim 1$. When the switching rate $\nu$ is increased further, the pdf becomes unimodal and sharply peaked about a value approaching $\mathcal{K}$, and eventually $N\approx\mathcal{K}$, when $\nu\gg 1$. The actual $N$ quasi-stationary distribution shows slightly broader peaks, due to demographic noise, but is otherwise well described by the PDMP approximation [11, 12, 40, 41]. In particular, the average of $N$ over the $N$-PDMP pdf given by
@@
is an accurate approximation of $\langle N\rangle$, the actual stationary mean population size [11, 12, 40, 13], with $\langle N\rangle\approx\langle N\rangle_{\nu,\delta}\approx[(1+\delta)K_{+}+(1-\delta)K_{-}]/2$ when $\nu\ll 1$ and $\langle N\rangle\approx\langle N\rangle_{\nu,\delta}\approx\mathcal{K}$ when $\nu\gg 1$. As shown in references [11, 12, 40, 13], equation (16) correctly captures that, at fixed $\delta$, the average population size $\langle N\rangle$ is a decreasing function of $\nu$.
Figure 3: Eco-evolutionary dynamics sample paths. (a)-(d) Examples of $R$ fraction sample paths in silico (blue lines) for four concentration cooperation thresholds $x_{\text{th}}$ (dashed black lines), three average switching rates $\nu$, and two environmental biases $\delta$, for $s=0.1$, $a=0.25$, $K_{-}=100$, and $K_{+}=1000$, with initial conditions $N_{R}(t=0)=x_{\text{th}}K(t=0)$ and $N_{S}(t=0)=\left(1-x_{\text{th}}\right)K(t=0)$. Shaded and white areas in panels (b)-(c) encode periods of abundance ($K(t)=K_{+}$) and scarcity ($K_{-}$), respectively. Note the larger amplitude of demographic fluctuations for the latter since $N\rightarrow K_{-}\ll K_{+}$. (e)-(h) Same example paths as in the corresponding panels (a)-(d) for the population size $N$ (green lines); black lines show example paths for the carrying capacity $K$. The very high environmental switching rate $\nu$ in panel (h) provides the effectively constant carrying capacity $\mathcal{K}$ (black dashed line); see section 4.1. (i)-(l) Same example paths as in the previous panels for the number of $R$ (blue lines) and $S$ (red lines). Dashed lines show the corresponding (meta-)stable coexistence equilibrium $N_{R}=x_{\text{th}}K$ (blue) and $N_{S}=(1-x_{\text{th}})K$ (red); see section 3. The demographic noise (DN) and the low value of $x_{\text{th}}=0.1$ in panels (a),(e),(i) lead to fast extinction of $R$, whereas the high threshold $x_{\text{th}}=0.9$ and DN in panels (d),(h),(l) lead to an early fixation of $R$ and extinction of $S$.
@@ 4.3 Theory for the fixation-coexistence diagrams in fluctuating environments
In this section we develop the analytical methods for fluctuating environments allowing us to predict when, in our simulations, one of the strains will fixate or go extinct, and when both strains will coexist for long-periods. The theory that we have devised and discuss here is used to analytically reproduce, see figure 4(e)-(h), the fixation-coexistence diagrams of figure 4(a)-(d) obtained from extensive computer simulations. Here, we show that microbial populations in time-varying environments can exhibit two qualitatively distinct behavioural regimes, depending on whether the environment has switched at least once. If this is the case, microbes experience a mixture of the two possible environments, and dynamics is effectively captured by averaging fixation/extinction probability fluxes over the total population distribution. When switches are unlikely to occur, the environment is essentially static and the population can be assumed to coincide with the initial value of the carrying capacity.
@@
We first define long-lived coexistence in fluctuating environments by direct extension of what has been discussed in static environments; see section 3.2. For given EV defined by the environmental mean switching rate and bias parameters $\{\nu,\delta\}$, there is long-lived coexistence of the strain $R$ and $S$ if the fixation time, here denoted by $\tau$, exceeds $2\langle N\rangle$ [32, 13]. Note that the mean fixation/extinction time, denoted by $\langle\tau\rangle$, coincides with the mean coexistence time, as in the environment-static case of section 3.2. We will thus say that, for given $\{\nu,\delta\}$, there is long-lived coexistence in the population of fluctuating size and volume, if $\tau>2\langle N\rangle\approx 2\langle N\rangle_{\nu,\delta}$. The rationale for this definition is the same as in the case of static environment, see section 3.2 and [89, 90, 91, 32, 13], with the difference that the constant population is now replaced by the average stationary population size $\langle N\rangle$ that we analytically approximate by $\langle N\rangle_{\nu,\delta}$ obtained from the $N$-PDMP according to equation (16). Therefore, we will say that, for given environmental parameters $\{\nu,\delta\}$, there is fast fixation or, equivalently, early extinction of a strain, if $R$ or $S$ fixates in a time $\tau$ that is less than $2\langle N\rangle\approx 2\langle N\rangle_{\nu,\delta}$.
@@
Secondly, we distinguish two fluctuating environment regimes in which the fixation and long-lived coexistence probability are evaluated by distinct suitable approaches. We refer to these different behaviours as the quenched (Q) and annealed (A) regimes. Dynamic environments in the regime Q are so slow, $\nu\ll 1$, and/or so biased, $\delta\approx\pm 1$, that the environment is unlikely to experience any switch by time $2\langle N\rangle_{\nu,\delta}$ regardless of the initial condition; whereas, in the same time, environments in regime A experience at least one environmental switch. Similar Q and A environmental regimes have been identified elsewhere, see e.g. [96, 97] [^4] It is worth noting that the annealed regime often corresponds to the limit $\nu\to\infty$ where the environmental noise $\xi$ self averages [96, 97, 11, 12, 40, 41]. Here, the case of intermediate switching with a finite number of switches prior to fixation/extinction is qualitatively similar to the case $\nu\to\infty$, and for simplicity we here refer to it as the annealed regime A.. The border between Q and A regimes is determined by
@@
which splits the fixation-coexistence $\nu-\delta$ diagrams of figure 4 in two regions separated by green/yellow lines: where $1/\left(\nu(1-|\delta|)\right)>2\langle N\rangle_{\nu,\delta}$, the average time for an environmental switch to occur exceeds the reference time for long-lived coexistence. This means that in this region $K$ is unlikely to experience any switch by time $2\langle N\rangle_{\nu,\delta}$, and the previous inequality hence delineates the Q regime. Here, the A regime is thus defined by $1/\left(\nu(1-|\delta|)\right)<2\langle N\rangle_{\nu,\delta}$, where, on average, $K$ switches at least once by time $2\langle N\rangle_{\nu,\delta}$.
Quenched environmental regime. In regime Q (on the left of green/yellow lines of figure 4), the carrying capacity remains at initial value, $K(t)=K(0)$, at least until time $2\langle N\rangle_{\nu,\delta}$. Since the environment is already in its quasi-stationary state, this initial value is either $K(0)=K_{+}$, with probability $(1+\delta)/2$, or $K(0)=K_{-}$, with probability $(1-\delta)/2$. Hence, the joint probability of $R$ fixating and that fixation/extinction occurs in a time $\tau<2\langle N\rangle_{\nu,\delta}$ (fast fixation) in the regime Q, denoted by $\Phi^{\text{Q}}$, is
@@
is the probability of long-lived coexistence in the Q regime, starting (and staying) at a total population $N$, with mean coexistence time (MCT) $\langle\tau_{N}\rangle$, where $\eta_{N}^{Q}$ is defined as the environmental-dynamic counterpart of the environmental-static equation (12).
Annealed environmental regime. In regime A (on the right of green/yellow lines of figure 4), the environment switches at least once prior to time $2\langle N\rangle_{\nu,\delta}$ regardless of the initial condition. In this regime, $N$ therefore experiences a broad range of values in sampling its quasi-stationary distribution, here approximated by $p_{\nu,\delta}(N)$ of equation (15), before a fixation/extinction event or long-lived coexistence occurs. In the vein of [11, 12], we thus assume that when a fixation event occurs, it arises in a population of size $N$ ponderated by the probabilistic weight $p_{\nu,\delta}(N)$. Following [92, 87], we know that the MCT $\langle\tau_{N}\rangle$ in a population of constant size $N$ is the inverse of the probability flux towards the absorbing boundaries, and the fixation probability $\phi_{N}$ is given by the relative flux into the absorbing state $N_{R}=N$ [87]. Hence, $\phi_{N}/\langle\tau_{N}\rangle$ is defined as the “rate to $R$ fixation” in a population of size $N$, and similarly we define $(1-\phi_{N})/\langle\tau_{N}\rangle$ as the “rate to $R$ extinction” or, equivalently, as the “rate to $S$ fixation”. In regime A, we assume that any fixation/extinction events occur while $N$ is at quasi-stationarity and, by analogy with the $N$-constant case, experiences a rate to $R$ fixation/extinction $\phi_{N}/\langle\tau_{N}\rangle$ and $(1-\phi_{N})/\langle\tau_{N}\rangle$, respectively; but we now average these over $p_{\nu,\delta}(N)$ to account for the varying population size in fluctuating environments. We thus obtain the effective rates of $R$ fixation/extinction in regime A, respectively given by
@@
Since $\langle\tau_{K_{+}}\rangle\gg\langle\tau_{K_{-}}\rangle$, the main contributions to $\mathcal{T}_{F/E}$, and hence to fixation/extinction, arise from $N\approx K_{-}$. With $\mathcal{T}_{F/E}$, we obtain the $R$ fixation probability and mean coexistence time, $\phi^{\text{A}}$ and $\langle\tau^{\text{A}}\rangle$, respectively, in formal analogy with the environmental-static case [92, 87]:
@@
Moreover, for a dynamic environment of mean switching rate and bias parameters $\{\nu,\delta\}$, following the reasoning leading to equation (12), the long-lived coexistence probability in regime A is
@@
Crossover regime and general results. Having obtained the fixation and coexistence probabilities in regimes Q and A, we can suitably superpose their expressions to obtain predictions applicable in the crossover regime (about the green/yellow line in figure 4), as well as in regimes Q and A. This provides us with general results, that are valid for then entire range of environmental parameters $\{\nu,\delta\}$. Since the probability that no switches occur by time $2\langle N\rangle_{\nu,\delta}$ (regime Q) is
@@
and the probability that at least one switch has occurred by $2\langle N\rangle_{\nu,\delta}$ (regime A) is $1-\Pi$, the overall joint probability of $R$ fixation and fast fixation is
@@
The fixation-coexistence diagrams in figure 4(a)-(d) show simulation results for the probability of fast $R$ fixation (blue), fast $S$ fixation (red), or long-lived coexistence (black), see appendix section A, for different values of the concentration cooperation threshold $x_{\text{th}}$ when $N$ is at quasi-stationarity. The main features of these diagrams can be understood in terms of the analysis carried in section 3.2 and by referring to the critical cooperation threshold value $x_{\text{th}}^{*}\approx[\ln\left(1-s\right)]/[\ln\left(1-a\right)]$  [^5] In principle, a suitable environment-dynamic counterpart of $x_{\text{th}}^{*}$, given by equation (10), in a population of average size $\langle N\rangle_{\nu,\delta}$ is $x_{\text{th},\nu,\delta}^{*}\equiv\frac{\ln\left(1-s\right)}{\ln\left(1-a\right)}+\frac{1}{\langle N\rangle_{\nu,\delta}}\frac{\ln\left(\frac{s(1-a)}{a-s}\right)}{\ln\left(1-a\right)}$. However, when $\langle N\rangle\gg 1$ and $a$ is not too close to 1, as in our examples, $x_{\text{th},\nu,\delta}^{*}$ is well approximated by the leading contribution to $x_{\text{th}}^{*}$, and hence we can consider $x_{\text{th},\nu,\delta}^{*}\approx x_{\text{th}}^{*}\approx\frac{\ln\left(1-s\right)}{\ln\left(1-a\right)}$.: when $x_{\text{th}}<x_{\text{th}}^{*}$ (and $x_{\text{th}}\lesssim 1/\sqrt{\langle N\rangle_{\nu,\delta}}$ for significant demographic fluctuations), the most likely outcome is the fast fixation of $S$ (i.e., early extinction of $R$); while fast fixation of $R$ (early extinction of $S$) is the most probable outcome when $x_{\text{th}}>x_{\text{th}}^{*}$ (and $1-x_{\text{th}}\lesssim 1/\sqrt{\langle N\rangle_{\nu,\delta}}$); and long-lived coexistence of $R$ and $S$ is expected when $x_{\text{th}}$ is close to $x_{\text{th}}^{*}$ (and/or when ${\rm min}(x_{\text{th}},1-x_{\text{th}})\gg 1/\sqrt{\langle N\rangle_{\nu,\delta}}$ so that demographic fluctuations are negligible). For the parameters of figure 4, $x_{\text{th}}^{*}\approx 0.366$, and we indeed notice that most of figure 4(a) appears in red, panels (c) and (d) of figure 4 are mostly coded in blue, and most of figure 4(b) appears in black.